OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form

OpenStudy (anonymous):

\[y = 2x + c\]\[2(-1) + c = 5\]\[c = 5 + 2 = 7\]\[y = 2x + 7\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

A. 2x – y = –7 B.–2x – y = –7 C. 2x – y = 11 D. –2x + y = –11 this are my opinion :/

OpenStudy (anonymous):

Hang on standard form would be y - 2x - 7 = 0, right?

OpenStudy (anonymous):

(Which is equivalent to A)